### v1

# HybridNorm: Towards Stable and Efficient Transformer Training via Hybrid Normalization

Abstract
Transformers have become the de facto architecture for a wide range of machine learning tasks, particularly in large language models (LLMs). Despite their remarkable performance, challenges remain in training deep transformer networks, especially regarding the location of layer normalization. While Pre-Norm structures facilitate easier training due to their more prominent identity path, they often yield suboptimal performance compared to Post-Norm. In this paper, we propose HybridNorm, a straightforward yet effective hybrid normalization strategy that integrates the advantages of both Pre-Norm and Post-Norm approaches. Specifically, HybridNorm employs QKV normalization within the attention mechanism and Post-Norm in the feed-forward network (FFN) of each transformer block. This design not only stabilizes training but also enhances performance, particularly in the context of LLMs. Comprehensive experiments in both dense and sparse architectures show that HybridNorm consistently outperforms both Pre-Norm and Post-Norm approaches, achieving state-of-the-art results across various benchmarks. These findings highlight the potential of HybridNorm as a more stable and effective technique for improving the training and performance of deep transformer models. Code is available at https://github.com/BryceZhuo/HybridNorm.
Machine Learning, ICML
Figure 1: Training dynamics for 1.2B dense models with Pre-Norm, HybridNorm and HybridNorm under 1T training tokens. We present the training loss, validation loss, and downstream performance on HellaSwag and ARC-Easy (0.8 EMA smoothed), demonstrating that HybridNorm achieves superior performance.
Figure 2: Illustrations of different transformer layer structures: (a) Post-Norm architecture; (b) Pre-Norm architecture; (c) Pre-Norm with QK-Norm architecture; (d) HybridNorm architecture.
1 Introduction
Transformers have become the backbone of large language models (LLMs) and a wide range of machine learning applications. These architectures are capable of modeling long-range dependencies through self-attention mechanisms, which have made them the preferred choice for a variety of tasks, including language modeling, machine translation, and image processing (Devlin et al., 2019; Brown et al., 2020; Dosovitskiy et al., 2021). However, as transformer models become deeper and more complex, ensuring stable training remains a significant challenge. One critical factor that influences training stability is the choice of normalization methods, which is crucial for mitigating issues such as internal covariate shift and gradient instability (Xiong et al., 2020). Effectively addressing these challenges is crucial for fully harnessing the potential of deep transformer models in large-scale applications.
In transformers, Layer Normalization (LayerNorm) (Ba et al., 2016) plays a central role in stabilizing training by normalizing the activations within each layer. The two predominant strategies for applying LayerNorm are Pre-Layer Normalization (Pre-Norm) and Post-Layer Normalization (Post-Norm), each with its respective benefits and trade-offs. In the Pre-Norm architecture, normalization is applied before the residual addition, resulting in a more prominent identity path that facilitates faster convergence and more stable gradients (Xiong et al., 2020). This design is particularly advantageous when training deep models, as it helps mitigate gradient-related issues that can arise during backpropagation. However, while Pre-Norm can stabilize training, it often leads to inferior final performance compared to Post-Norm (Xie et al., 2023). In contrast, Post-Norm applies normalization after the residual connection, resulting in stronger regularization effects, which contribute to improved model performance. This approach has been shown to improve the generalization ability of transformers, particularly in very deep networks (Wang et al., 2024).
Despite the benefits of each approach, there is an inherent trade-off between training stability and final model performance. Pre-Norm structures typically stabilize training but may underperform in terms of generalization, while Post-Norm architectures provide better performance but can be more difficult to train, especially in deep models. To reconcile these trade-offs, we propose a hybrid normalization method that applies QKV normalization in the attention mechanism and Post-Norm in the feed-forward network (FFN), which is named as HybridNorm. The QKV normalization in the attention mechanism stabilizes the flow of information between layers by normalizing the query, key, and value components, while Post-Norm in the FFN ensures effective depth in the deeper layers of the transformer.
Through extensive experiments on large-scale models, we validate the effectiveness of our approach. Our results show that the hybrid normalization method significantly outperforms both Pre-Norm and Post-Norm across multiple benchmarks, providing a stable training process and improved model performance. Specifically, our method achieves superior results in the context of LLMs, where the benefits of both normalization schemes are most pronounced. We believe that this hybrid approach offers a promising solution for enhancing the training stability and performance of deep transformer architectures, particularly in the rapidly evolving field of LLMs.
The main contributions of this paper can be summarized as follows.
  • We propose HybridNorm, a novel hybrid normalization structure that combines the advantages of Pre-Norm and Post-Norm, offering a simple yet effective solution to enhancing performance in large transformer models. Our method is designed to exploit the strengths of both normalization approaches, ensuring robust convergence during training and superior final performance.
  • We provide an empirical analysis of the proposed hybrid normalization method, demonstrating its potential benefits in terms of gradient flow stability, regularization effects, and model robustness. This analysis highlights how our approach addresses the key challenges posed by deep transformer architectures.
  • Through extensive experiments on large-scale models, we empirically validate the effectiveness of our approach. Our results show that hybrid normalization significantly outperforms both Pre-Norm and Post-Norm across a variety of tasks, leading to more stable training and improved model performance, particularly in the context of large language models (LLMs).
The remainder of this paper is organized as follows. Section 2 reviews related work on normalization techniques in transformers. Section 3 introduces our proposed HybridNorm method, detailing its design and theoretical motivation. Section 4 presents the experimental setup and evaluation results, demonstrating the effectiveness of HybridNorm across various large-scale transformer models. Section 5 concludes the paper.
2 Related Work
Architecture Modifications in Transformers. Recent efforts in transformer architecture modifications have sought to optimize both the computational efficiency and the expressiveness of the model. These efforts include changes to the attention mechanism and feed-forward networks all aimed at improving performance on a variety of tasks, ranging from language modeling to vision tasks (Vaswani et al., 2017; Dosovitskiy et al., 2021). For example, Multi-head Latent Attention (MLA) (DeepSeek-AI, 2024), Mixture of Experts (MoE) (Yuksel et al., 2012). While these modifications contribute to more efficient training, they also require careful integration with other components, such as normalization layers, to maintain model stability and performance.
Normalization Types in Transformers. Normalization layers are integral to the success of deep learning models, and transformers are no exception. The most commonly used normalization technique in transformers is LayerNorm (Ba et al., 2016), which normalizes the activations of each layer independently. However, alternative methods such as RMSNorm (Zhang & Sennrich, 2019), which normalizes using root mean square statistics, have been proposed as more effective alternatives in certain settings. These methods are designed to mitigate the challenges of internal covariate shift and gradient instability, which are critical for the success of large-scale transformer models.
Normalization Settings in Attention. For training stability, QK-Norm (Henry et al., 2020; Dehghani et al., 2023) modifies the standard attention mechanism by applying normalization directly to the query (Q) and key (K) components during attention computation. Building upon this, QKV-Norm (Menary et al., 2024; Rybakov et al., 2024) extends the approach by normalizing the Query (Q), Key (K), and Value (V) components. This comprehensive normalization ensures that all critical components of the attention mechanism are normalized, resulting in enhanced stability and improved performance.
Location of Normalization Layers. Another line of research focuses on the location of the normalization. The choice between Pre-Norm and Post-Norm architectures has been widely studied in the transformer literature (Vaswani et al., 2017; Klein et al., 2017; Wang et al., 2019). Pre-Norm, where normalization is applied before the residual connection, has been shown to be more stable in deep networks and accelerates convergence (Xiong et al., 2020). Although Post-Norm is more challenging to train, it tends to deliver better final performance by normalizing after the residual connection (Liu et al., 2020). DeepNorm (Wang et al., 2024) was proposed as a strategy to address training instability in deep transformers, which scales the residual connections by a carefully chosen factor to improve gradient flow and mitigate exploding or vanishing gradients. The method most similar to ours is Mix-LN (Li et al., 2025), which applies Post-Norm to the earlier layers and Pre-Norm to the deeper layers, achieving improved training stability and better performance. In contrast, our HybridNorm integrates Pre-Norm and Post-Norm within each transformer block.
3 Method
In this section, we first review the two predominant normalization strategies in transformer architectures: Post-Norm and Pre-Norm. We then introduce our proposed hybrid normalization method, HybridNorm, and provide its formal definition.
3.1 Preliminaries
Scaled Dot-Product Attention. The scaled dot-product attention computes the attention scores between the Query (Q) and Key (K) matrices, scaled by the square root of the key dimension $d_{k}$, and applies these scores to the Value (V) matrix. The formulation is expressed as
$\mathrm{attn}(Q,K,V)=\mathrm{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V,$ (1)
where $Q,K,V\in\mathbb{R}^{n\times d_{k}}$ represent the query, key, and value matrices respectively, and $n$ is the sequence length.
Multi-Head Attention. Multi-head attention (MHA) extends the scaled dot-product attention mechanism by splitting the query, key, and value matrices into $h$ heads, each of size $d_{k}=d/h$. Each head independently computes attention scores, and the outputs are concatenated and linearly projected to the original dimension,
$\mathrm{MHA}(X)=\mathrm{Concat}(\mathrm{head}_{1},\dots,\mathrm{head}_{h})W^{O},$ (2)
where $\mathrm{head}_{i}=\mathrm{attn}(Q_{i},K_{i},V_{i})$ for $i=1,2,\dots,h$, $\{\bullet_{i}\}_{i=1}^{h}=\mathrm{Split}(XW_{\bullet})$ for $\bullet\in\{Q,K,V\}$, and $W_{Q},W_{K},W_{V},W_{O}\in\mathbb{R}^{d\times d}$ are learnable parameters. By enabling the model to focus on different subspaces of the input representation, MHA enhances the transformer’s capacity to capture diverse patterns in the input sequence.
Feed-Forward Networks (FFN). In transformers, the Feed-Forward Network (FFN) is applied independently to each position of the input sequence.
3.2 Post-Norm and Pre-Norm
The transformer architecture is composed of a stack of L blocks, each consisting of two key components: MHA and FFN. Residual connections and normalization layers are applied around both the MHA and FFN in each block to facilitate effective training and improve model stability. Figure 2 (a)&(b) illustrate Post-Norm and Pre-Norm, respectively.
Post-Norm. Post-Norm applies the normalization layer after the residual connection in each transformer sub-layer. Formally, the output of Post-Norm can be expressed as
$Y^{l}$ $=\mathrm{Norm}(\mathrm{MHA}(X^{l})+X^{l}),$ (3)
$X^{l+1}$ $=\mathrm{Norm}(\mathrm{FFN}(Y^{l})+Y^{l}),$ (4)
where $\mathrm{Norm}$ denotes RMSNorm (Zhang & Sennrich, 2019) or LayerNorm (Ba et al., 2016).
Pre-Norm. In contrast, Pre-Norm normalizes the input to the sub-layer, which allows for a more prominent identity path. The output of Pre-Norm is given by
$Y^{l}$ $=\mathrm{MHA}(\mathrm{Norm}(X^{l}))+X^{l},$ (5)
$X^{l+1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+Y^{l}.$ (6)
This structure facilitates better gradient flow and stable convergence, particularly for deep models. However, its reliance on normalization before the residual connection can lead to suboptimal performance compared to Post-Norm, as the normalization does not account for the interaction between the residual connection and the sub-layers output.
3.3 HybridNorm
To address the trade-offs between Post-Norm and Pre-Norm, we propose HybridNorm, a hybrid normalization strategy that integrates their strengths. Specifically, HybridNorm combines QKV-Norm (Menary et al., 2024; Rybakov et al., 2024) in MHA and Post-Norm in FFN.
QKV Normalization in Attention. In the attention mechanism, the query, key, and value matrices are normalized individually before computing the attention output. The normalized QKV matrices are then used in the scaled dot-product attention. QKV-Norm enhances the stability of model training and leads to improved downstream performance. Formally, attention with QKV-Norm is defined as
$\mathrm{attn}_{QKV}(Q,K,V)$ (7)
$=\mathrm{softmax}\left(\frac{\mathrm{Norm}(Q)\mathrm{Norm}(K)^{\top}}{\sqrt{d_{k}}}\right)\mathrm{Norm}(V).$
And we denote the multi-head attention with $\mathrm{attn}_{QKV}$ as $\mathrm{MHA}_{QKV}$.
HybridNorm Architecture. Combining the above, the overall output of a transformer block with HybridNorm can be expressed as
$Y^{l}$ $=\mathrm{MHA}_{QKV}(X^{l})+X^{l},$ (8)
$X^{l+1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+\mathrm{Norm}(Y^{l}).$ (9)
The architecture illustration can be found in Figure 2(d) and the pseudocode is shown in Algorithm 1. By integrating QKV normalization in the attention mechanism and Post-Norm in the FFN, HybridNorm achieves stable training dynamics and enhanced final performance. The theoretical gradient analysis can be found in Appendix A.
Remark 3.1.
The method most closely related to ours is Mix-LN (Li et al., 2025), which applies Post-Norm to the earlier layers and Pre-Norm to the deeper layers, resulting in enhanced training stability and performance. In contrast, our proposed HybridNorm integrates Pre-Norm and Post-Norm within each transformer block, providing a unified approach to leverage the benefits of both normalization strategies. Moreover, experiments demonstrate that HybridNorm achieves superior downstream performance compared to Mix-LN (see Table 4).
Special Treatment of First Block. Inspired by prior work (DeepSeek-AI, 2024), which employs the Mixture of Experts (MoE) architecture with specialized handling of the first layer, we explore the impact of introducing specialized normalization to the first transformer block. In our approach, the first layer of the transformer is treated differently by applying Pre-Norm on MHA and FFN, while maintaining QKV-Norm. Specifically, the structure of our first layer is defined as
$Y^{0}$ $=\mathrm{MHA}_{QKV}(\mathrm{Norm}(X^{0}))+X^{0},$ (10)
$X^{1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{0}))+Y^{0}.$ (11)
We refer to this variation of HybridNorm, which incorporates the specialized first block treatment, as HybridNorm. This design aims to stabilize the training of the first transformer block and boost overall performance by improving the flow of gradients in the early stages of training.
4 Experiments
Table 1: Downstream evaluation results of 1.2B dense models with Pre-Norm, HybridNorm, and HybridNorm under 1T training tokens.
Methods BasicArithmetic HellaSwag SciQ ARC-C ARC-E PIQA OpenbookQA COPA Avg.$\uparrow$
Pre-Norm 44.10 63.41 91.80 39.20 69.82 75.19 38.40 82.00 62.99
HybridNorm 44.12 64.22 91.88 39.13 71.05 74.72 38.88 82.00 63.25
HybridNorm 47.21 65.12 91.38 37.06 71.79 75.72 39.16 85.78 64.15
In this section, we demonstrate the effectiveness of HybridNorm through extensive experiments in LLMs.
4.1 Experiment Settings
Baseline. We evaluate HybridNorm across two series of models: dense models and Mixture of Experts (MoE) models. The dense models include two scales: 550M and 1B, with the latter containing approximately 1.27 billion parameters and utilizing an architecture similar to Llama 3.2 (Dubey et al., 2024). All analytical experiments are conducted on the 550M dense models. For the MoE model, we use the OLMoE framework (Muennighoff et al., 2025), which activates 1.3B parameters out of a total of 6.9B parameters (MoE-1B-7B). Both models are trained from scratch on the OLMoE Mix dataset  [^1] OLMoE Mix is available at https://huggingface.co/datasets/allenai/OLMoE-mix-0924. (Muennighoff et al., 2025).
Model Configuration. The 550M dense model has a model dimension of 1536, an FFN dimension of 4096, and utilizes 16 attention heads with 4 key/value heads per attention head. The 1.2B model features a larger model dimension of 2048 and an FFN dimension of 9192, with 32 attention heads and 8 key/value heads per attention head. The MoE-1B-7B model employs 16 attention heads, a model dimension of 2048, and an FFN dimension of 1024. Notably, it features 8 experts out of 64, providing a more fine-grained distribution of computational resources. All models consist of 16 layers and are trained with a consistent context length of 4096. More details can be found in Appendix B.
Hyperparameters. Model weights are initialized using Megatron initialization (Shoeybi et al., 2019) (See Section 4.4 for more details). For the optimization, we apply the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. All models are trained on sequences of 4096 tokens. For the dense model, we set the initial learning rate to 3e-4, decaying to 3e-5 using a cosine scheduler. The MoE model starts with a learning rate of 4e-4, decaying according to a cosine schedule. We summarize the hyperparameters in Table 7.
Evaluation Metrics. To evaluate the performance of LLMs with HybridNorm, we employ a diverse set of open benchmarks, including ARC-Easy (ARC-E) (Clark et al., 2018), ARC-Challenge (ARC-C) (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), CoQA (Reddy et al., 2019), Winogrande (Sakaguchi et al., 2021), MMLU (Hendrycks et al., 2021), BoolQ (Clark et al., 2019), COPA (Gordon et al., 2012), CSQA (Talmor et al., 2019), OBQA (Mihaylov et al., 2018), and SocialIQA (Sap et al., 2019). We leverage the LM Eval Harness (Gao et al., 2023) for standardized performance evaluation.
4.2 Main Results on Dense Models
We evaluate the performance of HybridNorm and HybridNorm on 1.2B dense transformer models. Figure 1 compares the training dynamics of dense models with different normalization methods. As shown in the figure, models with HybridNorm and HybridNorm exhibit consistently lower training loss and validation perplexity throughout training compared to Pre-Norm, highlighting their effectiveness in enhancing training stability and convergence.
Table 1 presents the downstream evaluation results across a diverse set of benchmarks. HybridNorm consistently outperforms Pre-Norm in most tasks, achieving the highest average score. Notably, it demonstrates substantial improvements in tasks such as BasicArithmetic (+3.11), HellaSwag (+1.71), and COPA (+3.78), indicating enhanced generalization and robustness. These results underscore the scalability of HybridNorm in larger transformer models, further validating its effectiveness in improving both training stability and downstream performance. More results can be found in Figure 7.
Figure 3: Training dynamics for MoE-1B-7B models with Pre-Norm and HybridNorm under 500B training tokens. We present the training loss, validation loss, and downstream performance on HellaSwag and MMLU Var, demonstrating that HybridNorm achieves superior performance.
Table 2: Downstream evaluation results of MoE-1B-7B with Pre-Norm and HybridNorm under 500B training tokens.
Methods HellaSwag ARC-C ARC-E PIQA WinoGrande OpenbookQA BoolQ COPA Avg.$\uparrow$
Pre-Norm 69.94 39.92 73.37 77.82 63.34 42.37 67.47 85.40 64.95
HybridNorm 70.71 42.27 75.77 78.06 64.58 43.18 68.41 86.00 66.12
4.3 Main Results on MoE Models
For MoE models, we conduct experiments on MoE-1B-7B with 8 experts selected from a pool of 64. Figure 3 presents the training dynamics of MoE models under different normalization strategies. Throughout the training, HybridNorm consistently achieves lower training loss and validation perplexity compared to Pre-Norm. These findings indicate that HybridNorm effectively alleviates optimization difficulties in large-scale MoE models, resulting in more stable training and enhanced downstream performance.
Further, as shown in Table 2, HybridNorm consistently outperforms Pre-Norm across various downstream tasks, achieving the highest average score. Notably, it demonstrates significant improvements in ARC-C (+2.35), ARC-E (+2.40), and OpenbookQA (+0.81), highlighting its ability to enhance generalization across diverse benchmarks.
4.4 Ablation Studies
Table 3: Training loss and validation perplexity of 550M dense models with Pre-Norm and HybridNorm under various initialization methods and 400B training tokens.
[…]
Pre-Norm Normal 2.75 20.29
Depth-Scaled 2.76 20.49
Megatron 2.76 20.44
HybridNorm Normal 2.76 20.44
Depth-Scaled 2.76 20.40
Megatron 2.74 20.00
Initialization. To evaluate the sensitivity of Pre-Norm and HybridNorm to initialization schemes, we conduct ablation studies comparing three widely used initialization strategies: Normal initialization (Nguyen & Salazar, 2019), Depth-Scaled initialization (Zhang et al., 2019; Gururangan et al., 2023), and Megatron initialization (Shoeybi et al., 2019). Normal initialization initializes all weights of linear layers using a truncated normal distribution with mean zero and standard deviation $1/\sqrt{2.5d}$, where $d$ is the hidden dimension. Depth-Scaled initialization and Megatron initialization introduce scaling factors to stabilize training in deep architectures. Specifically, Depth-Scaled initialization scales down the output projections of the attention and FFN by a factor of $\sqrt{2l}$, where $l$ is the layer index. In contrast, Megatron initialization scales down these projections by $\sqrt{2L}$, where $L$ is the total number of layers, mitigating gradient variance accumulation in very deep transformers.
As shown in Table 3, Pre-Norm and HybridNorm exhibit sensitivity across different initialization methods, achieving the lowest training loss and perplexity under Normal initialization and Megatron initialization, respectively. Therefore, we set the default initialization method for Pre-Norm to Normal initialization and for HybridNorm to Megatron initialization in all experiments, respectively, which ensures that even under settings that may be more favorable to baseline models, the superiority of our approach is effectively demonstrated.
Table 4: Abalation study of the position of normalization layers on 550M dense models with 400B tokens. We report the training loss, the perplexity on C4 and Pile, and the accuracy on HS (HellaSwag).
[…]
QKVC-Post 2.74 20.05 10.34 52.68
QKC-Post 2.73 20.00 10.31 52.26
QK-Post - diverge
KV-Post 2.74 20.11 10.38 52.10
KC-Post 2.75 20.34 10.47 51.15
Pre-QKV-Post 2.74 20.13 10.37 52.65
Pre-QKV-Pre 2.74 19.97 10.33 53.05
Pre-QK-Pre 2.75 20.22 10.43 52.29
QKV-Pre 2.74 19.96 10.33 52.57
Post-Norm 2.76 20.43 10.57 51.20
Pre-Norm 2.75 20.30 10.48 51.97
Mix-LN 2.76 20.43 10.56 51.29
Post-Pre 2.75 20.26 10.46 51.19
Pre-Post 2.74 20.15 10.40 52.42
HybridNorm 2.74 20.00 10.29 53.35
HybridNorm 2.73 19.85 10.25 53.36
Normalization Position. We investigate the impact of the position of normalization layers within the transformer block. First, we examine the effect of varying the placement of QKV normalization (e.g., normalization setting in attention). We extend the normalization setting by considering not only the Query (Q), Key (K), and Value (V) components but also the Context (C), which refers to the output of the attention mechanism. For instance, QKVC-Norm applies normalization to all four components: Query, Key, Value, and Context, while KV-Norm and KC-Norm focus on the normalization of the Key-Value and Key-Context pairs, respectively. QKVC-Post refers to transformer blocks that employ QKVC-Norm in the MHA while using Post-Norm in the FFN. Second, we explore the effect of integrating QKV-Norm into different transformer architectures. For instance, Pre-QKV-Pre refers to a configuration where QKV-Norm is applied with Pre-Norm in the MHA layer, while the FFN layer utilizes Post-Norm. Other configurations follow similar definitions. Finally, we compare various hybrid combinations of Pre-Norm and Post-Norm. Pre-Post refers to transformer blocks that apply Pre-Norm in the MHA and Post-Norm in the FFN, whereas Post-Pre adopts the opposite configuration. Mathematical formulas for methods mentioned above can be found in Appendix E.
As shown in Table 4, HybridNorm (also called QKV-Post) and its variant HybridNorm consistently surpass other configurations. Notably, HybridNorm achieves the lowest training loss and perplexity while attaining the highest accuracy on HellaSwag. Specifically, by comparing HybridNorm with the first block in Table 4, we find that QKV-Norm is the most effective normalization setting in the attention. Similarly, comparing HybridNorm with the second block, we observe that combining QKV-Norm with Post-Norm in the FFN yields superior performance. From the third block, one can see that the Pre-Post configuration indeed leads to improved performance, while replacing Pre-Norm in the MHA with QKV-Norm to form HybridNorm further enhances performance, achieving the best results.
Figure 4: Training loss and accuracy on HellaSwag of 550M dense models with different normalization methods for the first block.
Special Treatment of First Block. For the special treatment of the first block, we test different architectures, such as adding a normalization layer after embedding (call EmbedNorm) and armed the first block with QKV-norm and Pre-Norm in FNN (call First-QKV-Pre), which formulations are:
$Y^{0}$ $=\mathrm{MHA}_{QKV}(X^{0})+X^{0},$
$X^{1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{0}))+Y^{0}.$
As shown in Figure 4, we can see that, except for EmbedNorm, the special treatment of the first block effectively reduces training loss and improves downstream performance.
Figure 5: Gradient norm comparison of Pre-Norm, Post-Norm, and HybridNorm at different training steps.
Table 5: Performance of dense models with different depths under 400B training tokens. We report the training loss, the perplexity on C4 and Pile, and the accuracy on HellaSwag and PIQA.
Models 550M, 16 Layers 543M, 29 Layers
[…]
Post-Norm 2.76 20.43 10.57 51.20 71.80 diverge
Pre-Norm 2.75 20.30 10.48 51.97 71.14 2.73 19.88 10.31 53.86 71.63
HybridNorm 2.74 20.00 10.29 53.35 71.96 2.72 19.67 10.18 54.89 72.62
HybridNorm 2.73 19.85 10.25 53.36 71.15 2.71 19.52 10.10 54.54 72.01
4.5 Gradient Analysis
To gain deeper insights into the stability improvements introduced by HybridNorm, we analyze gradient norms across training iterations. As shown in Figure 5, we compare the gradient norms of Pre-Norm, Post-Norm, and HybridNorm at steps 1 and 100. The results indicate that Pre-Norm tends to exhibit gradient explosion in deeper models, while Post-Norm suffers from vanishing gradients, both of which hinder effective optimization. In contrast, HybridNorm maintains a well-balanced gradient flow throughout training, effectively mitigating these issues. An intuitive understanding is that Pre-Norm tends to amplify gradients, while Post-Norm diminishes them. HybridNorm alternates between these two normalization strategies, leading to more stable gradient propagation during backpropagation and effectively preventing gradient explosion or vanishing. This balanced gradient propagation contributes to smoother optimization dynamics and faster convergence, further reinforcing the effectiveness of HybridNorm in stabilizing transformer training.
Figure 6: Scaling law curves of Pre-Norm and HybridNorm.
4.6 Scaling Laws Experiments
We compare the loss scaling curves between Pre-Norm and HybridNorm across a range of dense model sizes, from 151M to 1.2B parameters. The model sizes used for the scaling law experiments are detailed in Table 6, and all models are trained using the same setting and hyperparameters for fair comparison, as specified in Table 7. Models with 151M, 285M, 550M, and 1.2B parameters are trained on 200B, 200B, 300B, and 1T tokens, respectively. As shown in Figure 6, HybridNorm exhibits superior scaling properties, demonstrating lower training loss as the model size increases. This highlights its capacity to maintain both training stability and performance, even for extremely large models, thereby making it highly suitable for scaling to billion-parameter regimes.
4.7 Deeper Models
To further evaluate the robustness of HybridNorm and HybridNorm in deeper architectures, we conduct experiments on transformers with depths ranging from 16 to 29 layers while maintaining a comparable parameter budget. This setup allows for a fair comparison of different normalization strategies in deep transformer architectures. As shown in Table 5, both HybridNorm and HybridNorm consistently outperform Pre-Norm and Post-Norm across various depths, demonstrating their effectiveness in stabilizing deep model training.
A particularly striking observation is that Post-Norm fails to converge at 29 layers, reinforcing its well-documented instability in deeper architectures. In contrast, HybridNorm and HybridNorm not only ensure stable training across all depths but also achieve significantly lower training loss and perplexity on both the C4 and Pile datasets. These improvements indicate that HybridNorm-based normalization strategies mitigate optimization difficulties that commonly arise in deep transformers. Furthermore, HybridNorm achieves the highest accuracy on challenging downstream benchmarks such as HellaSwag and PIQA, suggesting that its benefits extend beyond mere training stability to enhanced generalization on real-world tasks. These results provide strong empirical evidence that HybridNorm-based normalization schemes enable deeper transformer training while preserving superior optimization efficiency and downstream task performance.
5 Conclusion
In this paper, we have presented HybridNorm, a novel hybrid normalization approach that effectively integrates the strengths of Pre-Norm and Post-Norm to address the longstanding trade-offs in transformer training. We have provided empirical insights into how HybridNorm stabilizes gradient flow while maintaining strong regularization effects, demonstrating its potential to improve both convergence and final model performance. Extensive experiments across diverse benchmarks validate the efficacy of our approach, with results consistently showing that HybridNorm outperforms traditional normalization strategies in terms of stability and accuracy. Our findings underscore the necessity of re-evaluating normalization position within transformer architectures, thereby laying the groundwork for further exploration of hybrid strategies. We posit that HybridNorm signifies a substantial advancement in the design of more robust and performant deep transformer models, offering practical advantages for the training of next-generation large-scale neural networks.
Impact Statement
This paper proposes HybridNorm, a simple yet effective hybrid normalization technique that improves the training stability and performance of transformers. It has the potential to assist the LLM community in advancing transformer architectures and enhancing their overall effectiveness. While there may be societal implications of our work, none of which we feel must be specifically highlighted here.
Appendix A Theoretical Gradient Analysis
For simplicity, we consider a single-head attention layer. The input is $X\in{\mathbb{R}}^{s\times d}$, representing a sequence of $s$ tokens with dimension $d$. Throughout this section, we denote RMSNorm as $\mathrm{Norm}(\cdot)$, i.e., $\mathrm{Norm}({\bm{x}})=\alpha\odot\frac{x}{\mathrm{RMS}(x)}$ for ${\bm{x}}\in{\mathbb{R}}^{d}$, where $\mathrm{RMS}(x)=\sqrt{(x_{1}^{2}+\cdots+x_{d}^{2})/d}$. For further simplicity, we set $\alpha=\mathbf{1}_{d}$. The learnable parameters $W_{Q},W_{K},W_{V}\in{\mathbb{R}}^{d\times d_{k}}$ and $W_{O}\in{\mathbb{R}}^{d_{k}\times d}$. Let $X_{N}=\mathrm{Norm}(X)$, $M=\frac{1}{\sqrt{d_{k}}}X_{N}W_{Q}W_{K}^{\top}X_{N}^{\top}$, and $A=\mathrm{softmax}(M)$. The output of the attention block with Pre-Norm is then given by
$S=AX_{N}W_{V}W_{O}.$ (12)
Defining $Q=XW_{Q}$, $K=XW_{K}$, and $V=XW_{V}$, with their normalized counterparts $Q_{N}=\mathrm{Norm}(Q)$, $K_{N}=\mathrm{Norm}(K)$, and $V_{N}=\mathrm{Norm}(V)$, the output of the attention block with QKV-Norm is formulated as
$S_{N}=A_{N}V_{N}W_{O},$ (13)
here $A_{N}=\mathrm{softmax}(M_{N})$ and $M_{N}=\frac{1}{\sqrt{d_{k}}}Q_{N}K_{N}^{\top}$.
Following the prior work of Noci et al. (2022); Ormaniec et al. (2025), we analyze the gradients by computing derivatives using row-wise vectorization and arranging the Jacobian in the numerator layout, i.e.,
$\frac{\partial Y}{\partial X}=\frac{\partial\mathrm{vec}_{r}(Y)}{\partial\mathrm{vec}_{r}(X)^{\top}}.$
The following derivation primarily relies on the chain rule and the following rule
$\frac{\partial AWB}{\partial W}=A\otimes B^{\top},$ (14)
where $\otimes$ is the Kronecker product.
We first present the following extension of Lemma 2 in Noci et al. (2022).
Lemma A.1 ((Extention of Lemma 2 in Noci et al. (2022))).
The gradients of the attention with Pre-Norm defined in Eq. (12) are given by
$\frac{\partial S}{\partial W_{O}}$ $=\mathrm{softmax}\left(\frac{X_{N}W_{Q}W_{K}^{\top}X_{N}^{\top}}{\sqrt{d_{k}}}\right)X_{N}W_{V}\otimes I_{d},$ (15)
$\frac{\partial S}{\partial W_{V}}$ $=\mathrm{softmax}\left(\frac{X_{N}W_{Q}W_{K}^{\top}X_{N}^{\top}}{\sqrt{d_{k}}}\right)X_{N}\otimes W_{O}^{\top},$ (16)
$\frac{\partial S}{\partial W_{Q}}$ $=\left(I_{s}\otimes W_{O}^{\top}W_{V}^{\top}X_{N}^{\top}\right)\frac{\partial A}{\partial M}\left(\frac{X_{N}\otimes X_{N}W_{K}}{\sqrt{d_{k}}}\right),$ (17)
$\frac{\partial S}{\partial W_{K}}$ $=\left(I_{s}\otimes W_{O}^{\top}W_{V}^{\top}X_{N}^{\top}\right)\frac{\partial A}{\partial M}\left(\frac{X_{N}W_{Q}\otimes X_{N}}{\sqrt{d_{k}}}\right)K_{d_{k},d},$ (18)
where the gradients of the softmax with respect to its inputs is
$\frac{\partial A}{\partial M}=\mathrm{blockdiag}\left(\frac{\partial A_{i,:}}{\partial M_{i,:}}\right)=\mathrm{blockdiag}\left(\mathrm{diag}(A_{i,:})-A_{i,:}A_{i,:}^{\top}\right),$ (19)
$A_{i,:}$ is the $i$-th row of $A$ in column vector format, and the commutation matrix $K_{d_{k},d}$ is a permutation matrix that transforms the row-wise vectorization of $W_{K}$ into the column-wise vectorization of $W_{K}$, i.e.,
$K_{d_{k},d}\mathrm{vec}_{r}(W_{K})=\mathrm{vec}_{r}(W_{K}^{\top}).$
The gradient of $X_{N}$ with respect to $X$ is
$\frac{\partial X_{N}}{\partial X}=\frac{\partial\mathrm{Norm}(X)}{\partial X}=\mathrm{blockdiag}\left(\frac{\partial\mathrm{Norm}(X_{i,:})}{\partial X_{i,:}}\right)=\mathrm{blockdiag}\left(\frac{\sqrt{d}}{\|X_{i,:}\|_{2}}\left(I_{d}-\frac{X_{i,:}X_{i,:}^{\top}}{\|X_{i,:}\|_{2}^{2}}\right)\right),$ (20)
where $X_{i,:}$ is the $i$-th row of $X$ represented as a column vector. The definitions of $\frac{\partial Q_{N}}{\partial Q}$, $\frac{\partial K_{N}}{\partial K}$, and $\frac{\partial V_{N}}{\partial V}$ follow similarly, i.e., for $\bullet\in\{Q,K,V\}$,
$\frac{\partial\bullet_{N}}{\partial\bullet}=\mathrm{blockdiag}\left(\frac{\sqrt{d_{k}}}{\|\bullet_{i,:}\|_{2}}\left(I_{d_{k}}-\frac{\bullet_{i,:}\bullet_{i,:}^{\top}}{\|\bullet_{i,:}\|_{2}^{2}}\right)\right).$ (21)
Lemma A.2.
The gradients of the attention with QKV-Norm defined in Eq. (12) are
$\frac{\partial S_{N}}{\partial W_{O}}$ $=\mathrm{softmax}\left(\frac{\mathrm{Norm}(XW_{Q})\mathrm{Norm}(XW_{K})^{\top}}{\sqrt{d_{k}}}\right)\mathrm{Norm}(XW_{V})\otimes I_{d},$ (22)
$\frac{\partial S_{N}}{\partial W_{V}}$ $=\left(\mathrm{softmax}\left(\frac{\mathrm{Norm}(XW_{Q})\mathrm{Norm}(XW_{K})^{\top}}{\sqrt{d_{k}}}\right)\otimes W_{O}^{\top}\right)\frac{\partial V_{N}}{\partial V}(X\otimes I_{d_{k}}),$ (23)
$\frac{\partial S_{N}}{\partial W_{Q}}$ $=\left(I_{s}\otimes W_{O}^{\top}\mathrm{Norm}(XW_{V})^{\top}\right)\frac{\partial A_{N}}{\partial M_{N}}\left(\frac{I_{s}\otimes\mathrm{Norm}(XW_{K})}{\sqrt{d_{k}}}\right)\frac{\partial Q_{N}}{\partial Q}(X\otimes I_{d_{k}}),$ (24)
$\frac{\partial S_{N}}{\partial W_{K}}$ $=\left(I_{s}\otimes W_{O}^{\top}\mathrm{Norm}(XW_{V})^{\top}\right)\frac{\partial A_{N}}{\partial M_{N}}\left(\frac{\mathrm{Norm}(XW_{Q})\otimes I_{s}}{\sqrt{d_{k}}}\right)K_{d_{k},s}\frac{\partial K_{N}}{\partial K}(X\otimes I_{d_{k}}),$ (25)
where the definition of $\frac{\partial A_{N}}{\partial M_{N}}$ is similar to $\frac{\partial A}{\partial M}$ and $K_{d_{k},s}$ is the commutation matrix s.t., $K_{d_{k},s}\mathrm{vec}_{r}(K_{N})=\mathrm{vec}_{r}(K_{N}^{\top})$.
Proof.
For $\frac{\partial S_{N}}{\partial W_{O}}$, according to Eq. (14), we obtain
$\frac{\partial S_{N}}{\partial W_{O}}=A_{N}V_{N}\otimes I_{d}=\mathrm{softmax}\left(\frac{\mathrm{Norm}(XW_{Q})\mathrm{Norm}(XW_{K})^{\top}}{\sqrt{d_{k}}}\right)\mathrm{Norm}(XW_{V})\otimes I_{d}.$
For $\frac{\partial S_{N}}{\partial W_{V}}$, using the chain rule and Eq. (14), we have
$\frac{\partial S_{N}}{\partial W_{V}}$ $=\frac{\partial S_{N}}{\partial V_{N}}\frac{\partial V_{N}}{\partial V}\frac{\partial V}{\partial W_{V}}$
$=(A\otimes W_{O}^{\top})\frac{\partial V_{N}}{\partial V}(X\otimes I_{d_{k}})$
$=\left(\mathrm{softmax}\left(\frac{\mathrm{Norm}(XW_{Q})\mathrm{Norm}(XW_{K})^{\top}}{\sqrt{d_{k}}}\right)\otimes W_{O}^{\top}\right)\frac{\partial V_{N}}{\partial V}(X\otimes I_{d_{k}}).$
For $\frac{\partial S_{N}}{\partial W_{Q}}$, using the chain rule and Eq. (14), we obtain
$\frac{\partial S_{N}}{\partial W_{Q}}$ $=\frac{\partial S_{N}}{\partial A_{N}}\frac{\partial A_{N}}{\partial M_{N}}\frac{\partial M_{N}}{\partial Q_{N}}\frac{\partial Q_{N}}{\partial Q}\frac{\partial Q}{W_{Q}}$
$=\left(I_{s}\otimes W_{O}^{\top}V_{N}^{\top}\right)\frac{\partial A_{N}}{\partial M_{N}}\left(\frac{I_{s}\otimes K_{N}}{\sqrt{d_{k}}}\right)\frac{\partial Q_{N}}{\partial Q}(X\otimes I_{d_{k}})$
$=\left(I_{s}\otimes W_{O}^{\top}\mathrm{Norm}(XW_{V})^{\top}\right)\frac{\partial A_{N}}{\partial M_{N}}\left(\frac{I_{s}\otimes\mathrm{Norm}(XW_{K})}{d_{k}}\right)\frac{\partial Q_{N}}{\partial Q}(X\otimes I_{d_{k}}).$
Similarly, for $\frac{\partial S_{N}}{\partial W_{K}}$, we have
$\frac{\partial S_{N}}{\partial W_{K}}$ $=\frac{\partial S_{N}}{\partial A_{N}}\frac{\partial A_{N}}{\partial M_{N}}\frac{\partial M_{N}}{\partial K_{N}}\frac{\partial K_{N}}{\partial K}\frac{\partial K}{W_{K}}$
$=\left(I_{s}\otimes W_{O}^{\top}V_{N}^{\top}\right)\frac{\partial A_{N}}{\partial M_{N}}\left(\frac{Q_{N}\otimes I_{s}}{d_{k}}\right)\frac{\partial\mathrm{vec}_{r}(K_{N}^{\top})}{\partial\mathrm{vec}_{r}(K_{N})^{\top}}\frac{\partial K_{N}}{\partial K}(X\otimes I_{d_{k}})$
$=\left(I_{s}\otimes W_{O}^{\top}\mathrm{Norm}(XW_{V})^{\top}\right)\frac{\partial A_{N}}{\partial M_{N}}\left(\frac{\mathrm{Norm}(XW_{Q})\otimes I_{s}}{\sqrt{d_{k}}}\right)K_{d_{k},s}\frac{\partial K_{N}}{\partial K}(X\otimes I_{d_{k}}).$
Theorem A.3.
For the attention with Pre-Norm, we have
$\left\|\frac{\partial S}{\partial W_{O}}\right\|_{F}$ $=\mathcal{O}\left(sd\|W_{V}\|_{2}\right),$ (26)
$\left\|\frac{\partial S}{\partial W_{V}}\right\|_{F}$ $=\mathcal{O}\left(s\sqrt{d}\|W_{O}\|_{F}\right),$ (27)
$\left\|\frac{\partial S}{\partial W_{Q}}\right\|_{F}$ $=\mathcal{O}\left(\frac{(sd)^{3/2}}{2\sqrt{d_{k}}}\|W_{K}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2}\right),$ (28)
$\left\|\frac{\partial S}{\partial W_{K}}\right\|_{F}$ $=\mathcal{O}\left(\frac{(sd)^{3/2}}{\sqrt{d_{k}}}\|W_{Q}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2}\right).$ (29)
For the attention with QKV-Norm, we have
$\left\|\frac{\partial S_{N}}{\partial W_{O}}\right\|_{F}$ $=\mathcal{O}\left(sd\right),$ (30)
$\left\|\frac{\partial S_{N}}{\partial W_{V}}\right\|_{F}$ $=\mathcal{O}\left(\frac{s\sqrt{d_{k}}}{\sigma_{\min}^{V}}\|W_{O}\|_{2}\right),$ (31)
$\left\|\frac{\partial S_{N}}{\partial W_{Q}}\right\|_{F}$ $=\mathcal{O}\left(\frac{s^{3/2}d_{k}}{\sigma_{\min}^{Q}}\|W_{O}\|_{2}\right),$ (32)
$\left\|\frac{\partial S_{N}}{\partial W_{K}}\right\|_{F}$ $=\mathcal{O}\left(\frac{s^{3/2}d_{k}}{\sigma_{\min}^{K}}\|W_{O}\|_{2}\right),$ (33)
where $\sigma_{\min}^{Q},\sigma_{\min}^{K},\sigma_{\min}^{V}$ are minimal singular value of $XW_{Q},XW_{K},XW_{V}$, respectivaly.
Theorem A.3 presents the gradient norms of different methods. In the attention with Pre-Norm, the gradient of the weight matrix exhibits strong dependencies on other weights; for instance, $W_{Q}$ and $W_{K}$ are influenced by all three other weight matrices but not by themselves. In contrast, in the attention with QKV-Norm, the gradient of each weight matrix depends at most on itself and $W_{O}$. This suggests that the gradient of the attention with Pre-Norm is more tightly coupled with other weight matrices compared to the gradient of the attention with QKV-Norm. Therefore, during the gradient optimization process, if the norm of a certain weight becomes excessively large, it is more challenging to control in the attention with Pre-Norm, leading to an increase in gradient magnitude. This, in turn, creates a vicious cycle that may result in model collapse. In contrast, the attention with QKV-Norm alleviates this issue to some extent, which significantly benefits the stability of model training.
The proof is primarily based on the following facts
  • $\mathrm{tr}(B\otimes C)=\mathrm{tr}(B)\mathrm{tr}(C)$
  • $(B\otimes C)(D\otimes E)=(BD)\otimes(CE)$
  • $\|B\otimes C\|_{F}=\|B\|_{F}\|C\|_{F}$
  • $\|B\otimes C\|_{2}=\|B\|_{2}\|C\|_{2}$
  • $\|BC\|_{2}\leq\|B\|_{2}\|C\|_{2}$
  • $\|BC\|_{F}\leq\|B\|_{2}\|C\|_{F}\leq\|B\|_{F}\|C\|_{F}$
  • If ${\bm{p}}\in{\mathbb{R}}^{s}$, $p_{i}\geq 0$ and $\sum_{i=1}^{s}$, then $\|\mathrm{diag}({\bm{p}})-{\bm{p}}{\bm{p}}^{\top}\|_{2}\leq\frac{1}{2}$
  • If $A\in{\mathbb{R}}^{s\times s}$ is a stochastic matrix, i.e., $A\mathbf{1}_{s}=\mathbf{1}_{s}$ and each entry is nonnegative, then $\|A\|_{2}\leq\|A\|_{F}\leq\sqrt{s}$
  • $\|X_{N}\|_{F}=\sqrt{sd}$
Proof.
According to fundamental algebraic operations, we obtain
$\left\|\frac{\partial S}{\partial W_{O}}\right\|_{F}=\left\|AX_{N}W_{V}\otimes I_{d}\right\|_{F}=\|AX_{N}W_{V}\|_{F}\|I_{d}\|_{F}\leq\sqrt{d}\|A\|_{2}\|X_{N}\|_{F}\|W_{V}\|_{2}\leq sd\|W_{V}\|_{2},$
$\left\|\frac{\partial S}{\partial W_{V}}\right\|_{F}=\left\|AX_{N}\otimes W_{O}^{\top}\right\|_{F}=\|AX_{N}\|_{F}\|W_{O}\|_{F}\leq\|A\|_{2}\|X_{N}\|_{F}\|W_{O}\|_{F}\leq s\sqrt{d}\|W_{O}\|_{F},$
$\left\|\frac{\partial S}{\partial W_{Q}}\right\|_{F}$ $=\left\|\left(I_{s}\otimes W_{O}^{\top}W_{V}^{\top}X_{N}^{\top}\right)\frac{\partial A}{\partial M}\left(\frac{X_{N}\otimes X_{N}W_{K}}{\sqrt{d_{k}}}\right)\right\|_{F}$
$=\frac{1}{\sqrt{d_{k}}}\left\|\left(I_{s}\otimes W_{O}^{\top}W_{V}^{\top}X_{N}^{\top}\right)\mathrm{blockdiag}(\mathrm{diag}(A_{i,:})-A_{i,:}A_{i,:}^{\top})(X_{N}\otimes X_{N}W_{K})\right\|_{F}$
$=\frac{1}{\sqrt{d_{k}}}\left\|\mathrm{blockdiag}((W_{O}^{\top}W_{V}^{\top}X_{N}^{\top})(\mathrm{diag}(A_{i,:})-A_{i,:}A_{i,:}^{\top}))(X_{N}\otimes X_{N}W_{K})\right\|_{F}$
$=\frac{1}{\sqrt{d_{k}}}\left\|\left[\begin{array}[]{c}{X_{N}}_{1,:}^{\top}\otimes W_{O}^{\top}W_{V}^{\top}X_{N}^{\top}(\mathrm{diag}(A_{1,:})-A_{1,:}A_{1,:}^{\top}))X_{N}W_{K}\\ \vdots\\ {X_{N}}_{s,:}^{\top}\otimes W_{O}^{\top}W_{V}^{\top}X_{N}^{\top}(\mathrm{diag}(A_{s,:})-A_{s,:}A_{s,:}^{\top}))X_{N}W_{K}\end{array}\right]\right\|_{F}$
$\leq\frac{1}{\sqrt{d_{k}}}\|X_{N}\|_{F}\|W_{O}\|_{2}\|W_{V}\|_{2}\|X_{N}\|_{2}\frac{1}{2}\|W_{K}\|_{2}\|X_{N}\|_{F}$
$\leq\frac{1}{2\sqrt{d_{k}}}\|W_{K}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2}\|X_{N}\|_{F}^{3}$
$=\frac{(sd)^{3/2}}{2\sqrt{d_{k}}}\|W_{K}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2},$
$\left\|\frac{\partial S}{\partial W_{K}}\right\|_{F}$ $=\left\|\left(I_{s}\otimes W_{O}^{\top}W_{V}^{\top}X_{N}^{\top}\right)\frac{\partial A}{\partial M}\left(\frac{X_{N}W_{Q}\otimes X_{N}}{\sqrt{d_{k}}}\right)K_{d_{k},d}\right\|_{F}$
$=\left\|\left(I_{s}\otimes W_{O}^{\top}W_{V}^{\top}X_{N}^{\top}\right)\frac{\partial A}{\partial M}\left(\frac{X_{N}W_{Q}\otimes X_{N}}{\sqrt{d_{k}}}\right)\right\|_{F}$
$=\frac{1}{\sqrt{d_{k}}}\left\|\left(I_{s}\otimes W_{O}^{\top}W_{V}^{\top}X_{N}^{\top}\right)\mathrm{blockdiag}(\mathrm{diag}(A_{i,:})-A_{i,:}A_{i,:}^{\top})(X_{N}W_{Q}\otimes X_{N})\right\|_{F}$
$=\frac{1}{\sqrt{d_{k}}}\left\|\mathrm{blockdiag}((W_{O}^{\top}W_{V}^{\top}X_{N}^{\top})(\mathrm{diag}(A_{i,:})-A_{i,:}A_{i,:}^{\top}))(X_{N}W_{Q}\otimes X_{N})\right\|_{F}$
$=\frac{1}{\sqrt{d_{k}}}\left\|\left[\begin{array}[]{c}{(X_{N}W_{Q})}_{1,:}^{\top}\otimes W_{O}^{\top}W_{V}^{\top}X_{N}^{\top}(\mathrm{diag}(A_{1,:})-A_{1,:}A_{1,:}^{\top}))X_{N}\\ \vdots\\ (X_{N}W_{Q})_{s,:}^{\top}\otimes W_{O}^{\top}W_{V}^{\top}X_{N}^{\top}(\mathrm{diag}(A_{s,:})-A_{s,:}A_{s,:}^{\top}))X_{N}\end{array}\right]\right\|_{F}$
$\leq\frac{1}{\sqrt{d_{k}}}\|W_{Q}\|_{2}\|X_{N}\|_{F}\|W_{O}\|_{2}\|W_{V}\|_{2}\|X_{N}\|_{2}\frac{1}{2}\|X_{N}\|_{F}$
$\leq\frac{1}{2\sqrt{d_{k}}}\|W_{Q}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2}\|X_{N}\|_{F}^{3}$
$=\frac{(sd)^{3/2}}{2\sqrt{d_{k}}}\|W_{Q}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2}.$
Therefore,
$\left\|\frac{\partial S}{\partial W_{O}}\right\|_{F}$ $=\mathcal{O}\left(sd\|W_{V}\|_{2}\right),$
$\left\|\frac{\partial S}{\partial W_{V}}\right\|_{F}$ $=\mathcal{O}\left(s\sqrt{d}\|W_{O}\|_{F}\right),$
$\left\|\frac{\partial S}{\partial W_{Q}}\right\|_{F}$ $=\mathcal{O}\left(\frac{(sd)^{3/2}}{2\sqrt{d_{k}}}\|W_{K}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2}\right),$
$\left\|\frac{\partial S}{\partial W_{K}}\right\|_{F}$ $=\mathcal{O}\left(\frac{(sd)^{3/2}}{\sqrt{d_{k}}}\|W_{Q}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2}\right).$
As for the attention with QKV-Norm, we have
$\left\|\frac{\partial S_{N}}{\partial W_{O}}\right\|_{F}=\left\|A_{N}V_{N}\otimes I_{d}\right\|_{F}=\|A_{N}V_{N}\|_{F}\|I_{d}\|_{F}\leq\sqrt{d}\|A_{N}\|_{2}\|V_{N}\|_{F}\leq sd.$
For $\frac{\partial S_{N}}{\partial W_{V}}$, we have
$\left\|\frac{\partial S_{N}}{\partial W_{V}}\right\|_{F}$ $=\left\|(A\otimes W_{O}^{\top})\frac{\partial V_{N}}{\partial V}(X\otimes I_{d_{k}})\right\|_{F}$
$\leq\left\|A\otimes W_{O}^{\top}\right\|_{2}\left\|\frac{\partial V_{N}}{\partial V}(X\otimes I_{d_{k}})\right\|_{F}$
$\leq\sqrt{s}\|W_{O}\|_{2}\left\|\frac{\partial V_{N}}{\partial V}(X\otimes I_{d_{k}})\right\|_{F}.$
According to Eq. (21), we get
$\left\|\frac{\partial V_{N}}{\partial V}(X\otimes I_{d_{k}})\right\|_{F}$ $=\left\|\mathrm{blockdiag}\left(\frac{\sqrt{d_{k}}}{\|{V}_{i,:}\|_{2}}\left(I_{d_{k}}-\frac{{V}_{i,:}{V}_{i,:}^{\top}}{\|{V}_{i,:}\|_{2}^{2}}\right)\right)(X\otimes I_{d_{k}})\right\|_{F}$
$=\left\|\left[\begin{array}[]{c}{X}_{1,:}^{\top}\otimes\left(\frac{\sqrt{d_{k}}}{\|{V}_{1,:}\|_{2}}\left(I_{d_{k}}-\frac{{V}_{1,:}{V}_{1,:}^{\top}}{\|{V}_{1,:}\|_{2}^{2}}\right)\right)\\ \vdots\\ X_{s,:}^{\top}\otimes\left(\frac{\sqrt{d_{k}}}{\|{V}_{s,:}\|_{2}}\left(I_{d_{k}}-\frac{{V}_{s,:}{V}_{s,:}^{\top}}{\|{V}_{s,:}\|_{2}^{2}}\right)\right)\end{array}\right]\right\|_{F}$
$\leq\sqrt{d_{k}\sum_{i=1}^{s}\frac{\|X_{i,:}\|_{2}^{2}}{\|V_{i,:}\|_{2}^{2}}}$
$=\sqrt{d_{k}\sum_{i=1}^{s}\frac{\|X_{i,:}\|_{2}^{2}}{\|W_{V}^{\top}X_{i,:}\|_{2}^{2}}}$
$\leq\frac{\sqrt{sd_{k}}}{\sigma_{\min}^{V}}.$
Similarly, we can get
$\left\|\frac{\partial Q_{N}}{\partial Q}(X\otimes I_{d_{k}})\right\|_{F}\leq\frac{\sqrt{sd_{k}}}{\sigma_{\min}^{Q}},\quad\left\|\frac{\partial K_{N}}{\partial K}(X\otimes I_{d_{k}})\right\|_{F}\leq\frac{\sqrt{sd_{k}}}{\sigma_{\min}^{K}}.$
It follows that
$\left\|\frac{\partial S_{N}}{\partial W_{V}}\right\|_{F}\leq\frac{s\sqrt{d_{k}}}{\sigma_{\min}^{V}}\|W_{O}\|_{2}.$ (34)
For $\frac{\partial S_{N}}{\partial W_{Q}}$, we have
$\left\|\frac{\partial S_{N}}{\partial W_{Q}}\right\|_{F}$ $=\left\|\left(I_{s}\otimes W_{O}^{\top}V_{N}^{\top}\right)\frac{\partial A_{N}}{\partial M_{N}}\left(\frac{I_{s}\otimes K_{N}}{\sqrt{d_{k}}}\right)\frac{\partial Q_{N}}{\partial Q}(X\otimes I_{d_{k}})\right\|_{F}$
$\leq\frac{1}{\sqrt{d_{k}}}\left\|\left(I_{s}\otimes W_{O}^{\top}V_{N}^{\top}\right)\frac{\partial A_{N}}{\partial M_{N}}\left(I_{s}\otimes K_{N}\right)\right\|_{2}\left\|\frac{\partial Q_{N}}{\partial Q}(X\otimes I_{d_{k}})\right\|_{F}$
$\leq\frac{1}{\sqrt{d_{k}}}\left\|\left(I_{s}\otimes W_{O}^{\top}V_{N}^{\top}\right)\mathrm{blockdiag}(\mathrm{diag}((A_{N})_{i,:})-(A_{N})_{i,:}(A_{N})_{i,:}^{\top})\left(I_{s}\otimes K_{N}\right)\right\|_{2}\frac{\sqrt{sd_{k}}}{\sigma_{\min}^{Q}}$
$=\frac{\sqrt{s}}{\sigma_{\min}^{Q}}\left\|\mathrm{blockdiag}(W_{O}^{\top}V_{N}^{\top}(\mathrm{diag}((A_{N})_{i,:})-(A_{N})_{i,:}(A_{N})_{i,:}^{\top})K_{N})\right\|_{2}$
$\leq\frac{\sqrt{s}}{\sigma_{\min}^{Q}}\|W_{O}\|_{2}\|V_{N}\|_{2}\frac{1}{2}\|K_{N}\|_{2}$
$\leq\frac{s^{3/2}d_{k}}{2\sigma_{\min}^{Q}}\|W_{O}\|_{2}.$
Similarly, for $\frac{\partial S_{N}}{\partial W_{K}}$, we have
$\left\|\frac{\partial S_{N}}{\partial W_{K}}\right\|_{F}\leq\frac{s^{3/2}d_{k}}{2\sigma_{\min}^{K}}\|W_{O}\|_{2}.$
Therefore,
$\left\|\frac{\partial S_{N}}{\partial W_{O}}\right\|_{F}$ $=\mathcal{O}\left(sd\right),$
$\left\|\frac{\partial S_{N}}{\partial W_{V}}\right\|_{F}$ $=\mathcal{O}\left(\frac{s\sqrt{d_{k}}}{\sigma_{\min}^{V}}\|W_{O}\|_{2}\right),$
$\left\|\frac{\partial S_{N}}{\partial W_{Q}}\right\|_{F}$ $=\mathcal{O}\left(\frac{s^{3/2}d_{k}}{\sigma_{\min}^{Q}}\|W_{O}\|_{2}\right),$
$\left\|\frac{\partial S_{N}}{\partial W_{K}}\right\|_{F}$ $=\mathcal{O}\left(\frac{s^{3/2}d_{k}}{\sigma_{\min}^{K}}\|W_{O}\|_{2}\right).$
Appendix B Details of Experiments
B.1 Architectures of Different Models
For dense models, we adopt a decoder-only transformer architecture akin to Llama 3.2 (Dubey et al., 2024), with model sizes ranging from 151M to 1.2B parameters. For the MoE model, we follow the structure of OLMoE (Muennighoff et al., 2025). The specific architecture of models is summarized in Table 6.
Table 6: Model architecture for dense models and MoE models.
Dense-151M Dense-285M Dense-550M Dense-543M Dense-1.2B MoE-1B-7B
Model Dimension 768 1024 1536 1024 2048 2048
FFN Dimension 2048 4096 4096 4096 9192 1024
Attention heads 16 16 16 16 32 16
Key/Value Heads 4 4 4 4 8 16
Layers 12 12 16 29 16 16
Vocabulary Size 100278 100278 100278 100278 100278 50280
Weight Typing True True True True True False
Context Length 4096 4096 4096 4096 4096 4096
Expert Granularity - - - - - 8 in 64
B.2 Hyperparameters for Pretraining
For the training hyperparameters, we primarily adopt the configuration outlined in OLMo 2 (OLMo et al., 2024) and OLMoE (Muennighoff et al., 2025). The training hyperparameters for our models across different sizes are presented in Table 7. The model is trained using the AdamW optimizer with a learning rate (LR) of 3e-4 (4e-4 for MoE), which is scheduled to decay following a cosine function. The minimum LR is set to 3e-5 (5e-5 for MoE) to prevent excessively small updates in the later stages of training. A weight decay of 0.1 is applied to regularize the model and prevent overfitting. The AdamW optimizer employs $\beta_{1}=0.9$ and $\beta_{2}=0.95$ to control the first and second momentum estimates, respectively. Gradient clipping is utilized with a threshold of 1 to mitigate the impact of large gradients during optimization. The model’s training also incorporates a warmup phase with a total of 8,388,608,000 tokens (10,485,760,000 for MoE). The initialization of the model’s parameters follows a normal distribution with a standard deviation defined as $1/\sqrt{2.5d}$, where $d$ is the model dimension. Furthermore, the initialization is truncated at 3 standard deviations to ensure a more stable starting point for training. The RoPE (Rotary Position Embedding) parameter $\theta$ is set to 500,000 (10000 for MoE), controlling the scale of position encodings. Finally, the activation function used in the model is SwiGLU, which has been shown to outperform traditional activation functions in various tasks.
Table 7: Hyperparameters for Pretraining.
Dense Model MoE-1B-7B
Optimizer AdamW AdamW
Learning Rate (LR) 3e-4 4e-4
Minimum LR 3e-5 5e-5
LR Schedule cosine cosine
Weight Decay 0.1 0.1
$\beta_{1}$ 0.9 0.9
$\beta_{2}$ 0.95 0.95
Gradient Clipping 1 1
Batch Size 1024 1024
Warmup Tokens 8,388,608,000 10,485,760,000
Init Distribution Megatron Megatron
Init std $1/\sqrt{2.5d}$ $1/\sqrt{2.5d}$
Init Truncation 3 std 3 std
RoPE $\theta$ 500000 10000
Activation SwiGLU SwiGLU
Load Balancing Loss Weight - 0.01
Router z-loss Weight - 0.001
Appendix C Additional Experimental Results
C.1 Overall Results for Dense Mdoels
Overall results for the dense model are presented in Figure 7, depicting validation losses and downstream evaluations over 1T training tokens. The comparison includes models with Pre-Norm, HybridNorm, and HybridNorm. One can see that both HybridNorm and HybridNorm outperform Pre-Norm, with HybridNorm achieving the lowest training and validation losses while delivering the best downstream performance.
Figure 7: Overall loss and downstream evaluations for the 1.2B dense models with 1T training tokens.
C.2 Overall Results for MoE Mdoels
Overall results for the MoE model are presented in Figure 8, illustrating validation losses and downstream evaluations over 500 billion training tokens. The comparison focuses on models employing Pre-Norm and HybridNorm. As shown, HybridNorm consistently achieves lower training and validation losses while delivering superior downstream performance.
Figure 8: Overall loss and downstream evaluations for the MoE models with 500B training tokens.
Appendix D PyTorch Style Implementation of HybridNorm
Algorithm 1 PyTorch style pseudocode for a Transformer block with HybridNorm
# q_norm, k_norm, v_norm, ffn_norm are normalization layers
# attn_proj and attn_out are linear layers
# attn is the attention
# ffn is the feedforward network
def forward(x):
# Attention block
res = x # shape (b, s, d)
q, k, v = attn_proj(x).split((d, d, d), dim=-1) # shape (b, s, d)
q, k, v = q.view(b, s, h, dk) # dk = d / h, h is the number of attention heads
q, k, v = q_norm(q), k_norm(k), v_norm(v)
x = attn(q,k,v) # shape (b, s, d)
x = attn_out(x) + res
# FFN block
x = ffn_norm(x)
x = ffn(x) + x
return x
Appendix E Formulas for Different Positions of Normalization Layers
In this section, we present the mathematical formulations for various normalization techniques. We begin by introducing the normalization layer within the attention mechanism.
Vanilla scaled dot-product attention are show in Eq. 1, and attention with QKV-Norm is defined in Eq. 7. Similarly, attention with QK-Norm is defined as
$\mathrm{attn}_{QK}(Q,K,V)=\mathrm{softmax}\left(\frac{\mathrm{Norm}(Q)\mathrm{Norm}(K)^{\top}}{\sqrt{d_{k}}}\right)V.$ (35)
Attention with KV-Norm is defined as
$\mathrm{attn}_{KV}(Q,K,V)=\mathrm{softmax}\left(\frac{Q\mathrm{Norm}(K)^{\top}}{\sqrt{d_{k}}}\right)\mathrm{Norm}(V).$ (36)
As mentioned in Section 4.4, we extend traditional normalization approaches by considering not only the Query (Q), Key (K), and Value (V) components but also the Context (C), which refers to the output of the attention mechanism. And attention with QKVC-Norm is defined as
$\mathrm{attn}_{QKVC}(Q,K,V)=\mathrm{Norm}\left(\mathrm{softmax}\left(\frac{\mathrm{Norm}(Q)\mathrm{Norm}(K)^{\top}}{\sqrt{d_{k}}}\right)\mathrm{Norm}(V)\right).$ (37)
Attention with QKC-Norm is defined as
$\mathrm{attn}_{QKC}(Q,K,V)=\mathrm{Norm}\left(\mathrm{softmax}\left(\frac{\mathrm{Norm}(Q)\mathrm{Norm}(K)^{\top}}{\sqrt{d_{k}}}\right)V\right).$ (38)
Attention with KC-Norm is defined as
$\mathrm{attn}_{KC}(Q,K,V)=\mathrm{Norm}\left(\mathrm{softmax}\left(\frac{Q\mathrm{Norm}(K)^{\top}}{\sqrt{d_{k}}}\right)V\right).$ (39)
Then we denote MHA with $\mathrm{attn}_{\#}$ as $\mathrm{MHA}_{\#}$ for #$\#\in\{QKVC,QKV,QKC,QK,KV,KC\}$,
$\mathrm{MHA}(X)_{\#}=\mathrm{Concat}(\mathrm{head}_{1},\dots,\mathrm{head}_{h})W^{O},$ (40)
where $\mathrm{head}_{i}=\mathrm{attn}_{\#}(Q_{i},K_{i},V_{i})$ for $i=1,2,\dots,h$, $\{\bullet_{i}\}_{i=1}^{h}=\mathrm{Split}(XW_{\bullet})$ for $\bullet\in\{Q,K,V\}$, and $W_{Q},W_{K},W_{V},W_{O}\in\mathbb{R}^{d\times d}$ are learnable parameters.
With the aforementioned definitions in hand, we present the mathematical formulations for the methods discussed in the Ablation Study below (#$\#\in\{QKVC,QKV,QKC,QK,KV,KC\}$).
##\##-Post:
$Y^{l}$ $=\mathrm{MHA}_{\#}(X^{l})+X^{l},$ (41)
$X^{l+1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+\mathrm{Norm}(Y^{l}).$ (42)
##\##-Pre:
$Y^{l}$ $=\mathrm{MHA}_{\#}(X^{l})+X^{l},$ (43)
$X^{l+1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+Y^{l}.$ (44)
Pre-##\##-Post:
$Y^{l}$ $=\mathrm{MHA}_{\#}(\mathrm{Norm}(X^{l}))+X^{l},$ (45)
$X^{l+1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+\mathrm{Norm}(Y^{l}).$ (46)
Pre-##\##-Pre:
$Y^{l}$ $=\mathrm{MHA}_{\#}(\mathrm{Norm}(X^{l}))+X^{l},$ (47)
$X^{l+1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+Y^{l}.$ (48)
Pre-Post:
$Y^{l}$ $=\mathrm{MHA}(\mathrm{Norm}(X^{l}))+X^{l},$ (49)
$X^{l+1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+\mathrm{Norm}(Y^{l}).$ (50)
Post-Pre:
$Y^{l}$ $=\mathrm{MHA}(\mathrm{Norm}(X^{l}))+\mathrm{Norm}(X^{l}),$ (51)
$X^{l+1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+Y^{l}.$ (52)