### v3

# HybridNorm: Towards Stable and Efficient Transformer Training via Hybrid Normalization

Abstract
Transformers have become the de facto architecture for a wide range of machine learning tasks, particularly in large language models (LLMs). Despite their remarkable performance, challenges remain in training deep transformer networks, especially regarding the position of layer normalization. While Pre-Norm structures facilitate more stable training owing to their stronger identity path, they often lead to suboptimal performance compared to Post-Norm. In this paper, we propose HybridNorm, a simple yet effective hybrid normalization strategy that integrates the advantages of both Pre-Norm and Post-Norm. Specifically, HybridNorm employs QKV normalization within the attention mechanism and Post-Norm in the feed-forward network (FFN) of each transformer block. We provide both theoretical insights and empirical evidence demonstrating that HybridNorm improves gradient flow and model robustness. Extensive experiments on large-scale transformer models, including both dense and sparse variants, show that HybridNorm consistently outperforms both Pre-Norm and Post-Norm approaches across multiple benchmarks. These findings highlight the potential of HybridNorm as a more stable and effective technique for improving the training and performance of deep transformer models.
[…]
1 Introduction
Transformers have become the backbone of large language models (LLMs) and a wide range of machine learning applications. These architectures are capable of modeling long-range dependencies through self-attention mechanisms, which have made them the preferred choice for a variety of tasks, including language modeling, machine translation, and image processing [8, 3, 10]. However, as transformer models become deeper and more complex, ensuring stable training remains a significant challenge. One critical factor that influences training stability is the choice of normalization methods, which is crucial for mitigating issues such as internal covariate shift and gradient instability [43]. Effectively addressing these challenges is crucial for fully harnessing the potential of deep transformer models in large-scale applications.
In transformers, Layer Normalization (LayerNorm) [1] plays a central role in stabilizing training by normalizing the activations within each layer. The two predominant strategies for applying LayerNorm are Pre-Layer Normalization (Pre-Norm) and Post-Layer Normalization (Post-Norm), each with its respective benefits and trade-offs. In the Pre-Norm architecture, normalization is applied before the residual addition, resulting in a more prominent identity path that facilitates faster convergence and more stable gradients [43]. This design is particularly advantageous when training deep models, as it helps mitigate gradient-related issues that can arise during backpropagation. However, while Pre-Norm can stabilize training, it often leads to inferior final performance compared to Post-Norm [42]. In contrast, Post-Norm applies normalization after the residual connection, resulting in stronger regularization effects, which contribute to improved model performance. This approach has been shown to improve the generalization ability of transformers, particularly in very deep networks [39]. Further discussion of related work is provided in Appendix 7..
Despite the benefits of each approach, there is an inherent trade-off between training stability and final model performance. Pre-Norm structures typically stabilize training but may underperform in terms of generalization, while Post-Norm architectures provide better performance but can be more difficult to train, especially in deep models. To reconcile these trade-offs, we propose a hybrid normalization method that applies QKV normalization in the attention mechanism and Post-Norm in the feed-forward network (FFN), which is named as HybridNorm. The QKV normalization in the attention mechanism stabilizes the flow of information between layers by normalizing the query, key, and value components, while Post-Norm in the FFN ensures the effective depth of the transformer.
Through extensive experiments on large-scale models, we validate the effectiveness of our approach. Our results show that the hybrid normalization method significantly outperforms both Pre-Norm and Post-Norm across multiple benchmarks, providing a stable training process and improved model performance. We believe that this hybrid approach offers a promising solution for enhancing the training stability and performance of deep transformer architectures, particularly in the rapidly evolving field of LLMs. The main contributions of this paper can be summarized as follows:
  • We propose HybridNorm, a novel hybrid normalization structure that combines the advantages of Pre-Norm and Post-Norm, offering a simple yet effective solution to enhancing performance in large transformer models. Our method is designed to exploit the strengths of both normalization approaches, ensuring robust convergence during training and superior final performance.
  • We present both theoretical and empirical analyses of HybridNorm, demonstrating its advantages in enhancing gradient flow stability and improving model robustness. Our findings underscore the method’s effectiveness in mitigating core challenges inherent to deep transformer architectures.
  • Through extensive experiments on large-scale models, we empirically validate the effectiveness of our approach. Our results show that hybrid normalization significantly outperforms both Pre-Norm and Post-Norm across a variety of tasks, leading to more stable training and improved model performance, particularly in the context of LLMs.
2 Preliminaries
Figure 1: Illustrations of different transformer layer structures: (a) Post-Norm architecture; (b) Pre-Norm architecture; (c) Pre-Norm with QK-Norm architecture; (d) HybridNorm architecture.
Scaled Dot-Product Attention
The scaled dot-product attention computes the attention scores between the Query (Q) and Key (K) matrices, scaled by the square root of the key dimension $d_{k}$, and applies these scores to the Value (V) matrix. The formulation is expressed as
$\mathrm{attn}(Q,K,V)=\mathrm{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V,$ (1)
where $Q,K,V\in\mathbb{R}^{n\times d_{k}}$ represent the query, key, and value matrices respectively, and $n$ is the sequence length.
Multi-Head Attention
Multi-head attention (MHA) extends the scaled dot-product attention mechanism by splitting the query, key, and value matrices into $h$ heads, each of size $d_{k}=d/h$. Each head independently computes attention scores, and the outputs are concatenated and linearly projected to the original dimension,
$\mathrm{MHA}(X)=\mathrm{Concat}(\mathrm{head}_{1},\dots,\mathrm{head}_{h})W^{O},$ (2)
where $\mathrm{head}_{i}=\mathrm{attn}(Q_{i},K_{i},V_{i})$ for $i=1,2,\dots,h$, $\{\bullet_{i}\}_{i=1}^{h}=\mathrm{Split}(XW_{\bullet})$ for $\bullet\in\{Q,K,V\}$, and learnable parameters $W_{Q},W_{K},W_{V},W_{O}\in\mathbb{R}^{d\times d}$ By enabling the model to focus on different subspaces of the input representation, MHA enhances the transformer’s capacity to capture diverse patterns in the input sequence.
2.1 Post-Norm and Pre-Norm
The transformer architecture is composed of a stack of L blocks, each consisting of two key components: MHA and FFN. Residual connections and normalization layers are applied around both the MHA and FFN in each block to facilitate effective training and improve model stability. Figure 1 (a)&(b) illustrate Post-Norm and Pre-Norm, respectively.
Post-Norm
Post-Norm applies the normalization layer after the residual connection in each transformer sub-layer. Formally, the output of Post-Norm can be expressed as
$Y^{l}$ $=\mathrm{Norm}(\mathrm{MHA}(X^{l})+X^{l}),\quad X^{l+1}=\mathrm{Norm}(\mathrm{FFN}(Y^{l})+Y^{l}),$ (3)
where $\mathrm{Norm}$ denotes RMSNorm [47] or LayerNorm [1].
Pre-Norm
In contrast, Pre-Norm normalizes the input to the sub-layer, which allows for a more prominent identity path. The output of Pre-Norm is given by
$Y^{l}$ $=\mathrm{MHA}(\mathrm{Norm}(X^{l}))+X^{l},\quad X^{l+1}=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+Y^{l}.$ (4)
This structure facilitates better gradient flow and stable convergence, particularly for deep models. However, its reliance on normalization before the residual connection can lead to suboptimal performance compared to Post-Norm, as the normalization does not account for the interaction between the residual connection and the sub-layers output. An analysis of the fundamental differences between the two approaches is provided in the Appendix 11.
3 HybridNorm
To address the trade-offs between Post-Norm and Pre-Norm, we propose HybridNorm, a hybrid normalization strategy that integrates their strengths. Specifically, HybridNorm combines QKV-Norm [22, 30] in MHA and Post-Norm in FFN.
QKV Normalization in Attention
In the attention mechanism, the query, key, and value matrices are normalized individually before computing the attention output. The normalized QKV matrices are then used in the scaled dot-product attention. QKV-Norm enhances the stability of model training and leads to improved downstream performance. Formally, attention with QKV-Norm is defined as
$\mathrm{attn}_{QKV}(Q,K,V)=\mathrm{softmax}\left(\frac{\mathrm{Norm}(Q)\mathrm{Norm}(K)^{\top}}{\sqrt{d_{k}}}\right)\mathrm{Norm}(V).$ (5)
And we denote the multi-head attention with $\mathrm{attn}_{QKV}$ as $\mathrm{MHA}_{QKV}$.
HybridNorm Architecture
Combining the above, the overall output of a transformer block with HybridNorm can be expressed as
$Y^{l}$ $=\mathrm{MHA}_{QKV}(X^{l})+X^{l},\quad X^{l+1}=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+\mathrm{Norm}(Y^{l}).$ (6)
The architecture illustration can be found in Figure 1(d) and the pseudocode is shown in Algorithm 1. By integrating QKV normalization in the attention mechanism and Post-Norm in the FFN, HybridNorm achieves stable training dynamics and enhanced final performance. The theoretical gradient analysis can be found in Appendix 8.
Remark 1.
The method most closely related to ours is Mix-LN [20], which applies Post-Norm to the earlier layers and Pre-Norm to the deeper layers, resulting in enhanced training stability and performance. In contrast, our proposed HybridNorm integrates Pre-Norm and Post-Norm within each transformer layer, thereby providing a more uniform approach across different layers to leverage the benefits of both normalization strategies. Moreover, experiments demonstrate that HybridNorm achieves superior downstream performance compared to Mix-LN (see Table 4 & Table 8).
Special Treatment of First Block
Inspired by prior work [6], which employs the Mixture of Experts (MoE) architecture with specialized handling of the first layer, we explore the impact of introducing specialized normalization to the first transformer block. In our approach, the first layer of the transformer is treated differently by applying Pre-Norm on MHA and FFN, while maintaining QKV-Norm. Specifically, the structure of our first layer is defined as
$Y^{0}$ $=\mathrm{MHA}_{QKV}(\mathrm{Norm}(X^{0}))+X^{0},\quad X^{1}=\mathrm{FFN}(\mathrm{Norm}(Y^{0}))+Y^{0}.$ (7)
We refer to this variation of HybridNorm, which incorporates the specialized first block treatment, as HybridNorm. This design aims to stabilize the training of the first transformer block and boost overall performance by improving the flow of gradients in the early stages of training.
(a) Layer gradient norm at step 1.
(b) Layer gradient norm at step 100.
Figure 2: Gradient norm of Pre-Norm, Post-Norm, and HybridNorm at different training steps.
4 Theoretical Analysis
4.1 Benefits of Hybrid Method
To gain deeper insights into the stability introduced by HybridNorm, we follow the approach of [35, 20] and analyze the evolution of gradient norms throughout training iterations. Suppose $x$ and $F$ are the input and the sublayer of Transformer, respectively. The output of Post-Norm is $y_{Post}=\mathrm{Norm}(x+F(x))$ and the output of Pre-Norm is $y_{Pre}=x+F(\mathrm{Norm}(x))$. Then we have
$\frac{\partial y_{Post}}{\partial x}$ $=\frac{\partial\mathrm{Norm}(x+F(x))}{\partial(x+F(x))}\left(I+\frac{\partial F(x)}{\partial x}\right),$ (8)
$\frac{\partial y_{Pre}}{\partial x}$ $=I+\frac{\partial F(\mathrm{Norm}(x))}{\partial\mathrm{Norm}(x)}\frac{\partial\mathrm{Norm}(x)}{\partial x},$ (9)
where the gradient of normalization is $\frac{\partial\mathrm{Norm}(x)}{\partial x}=\alpha\odot\frac{\sqrt{d}}{||x||_{2}}\left(I-\frac{xx^{\top}}{||x||_{2}^{2}}\right)$. The gradient of Post-Norm is the product of two gradients, one of which is the normalization. If the spectral radius of the normalization gradient is less than 1, it causes gradient vanishing. In Pre-Norm, the residual connection is isolated from the normalization, ensuring that gradients retain a lower bound, thereby preventing gradient vanishing. To ensure relatively stable gradients, a natural idea is to use both types of normalization within a Transformer block, leading to Pre-Post (Pre-Norm in MHA and Post-Norm in FFN) and Post-Pre. From Table 4, we find that Pre-Post achieves the best performance, which is why we adopt this. And placing Pre-Norm in MHA with QKV-Norm further enhances performance.
In Figure 2, we compare the gradient norms of Pre-Norm, Post-Norm, and HybridNorm at steps 1 and 100. The results indicate that Pre-Norm tends to exhibit gradient explosion in deeper models, while Post-Norm suffers from vanishing gradients, both of which hinder effective optimization. In contrast, HybridNorm maintains a well-balanced gradient flow throughout training, effectively mitigating these issues. An intuitive understanding is that Pre-Norm tends to amplify gradients, while Post-Norm diminishes them. HybridNorm alternates between these two normalization strategies, leading to more stable gradient propagation during backpropagation and effectively preventing gradient explosion or vanishing. This balanced gradient propagation contributes to smoother optimization dynamics and faster convergence, further reinforcing the effectiveness of HybridNorm in stabilizing training.
4.2 Benefits of QKV-Norm
Theoretically, we study how QKV-Norm affects the gradient flow during backpropagation, which is crucial for training stability in deep transformer models. Our analysis reveals that QKV-Norm helps decouple gradients between different weight matrices, thereby stabilizing training.
Theorem 1 (Informal version of Theorem 2).
Suppose the the output of the attention is $S$, the input $X\in\mathbb{R}^{s\times d}$, parameters $W_{Q},W_{K},W_{V},W_{O}^{\top}\in\mathbb{R}^{d\times d_{k}}$. For the attention with Pre-Norm, we have
$\left\|\frac{\partial S}{\partial W_{O}}\right\|_{F}$ $=\mathcal{O}\left(\|W_{V}\|_{2}\right),\left\|\frac{\partial S}{\partial W_{V}}\right\|_{F}=\mathcal{O}\left(\|W_{O}\|_{F}\right),$
$\left\|\frac{\partial S}{\partial W_{Q}}\right\|_{F}$ $=\mathcal{O}\left(\|W_{K}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2}\right),\left\|\frac{\partial S}{\partial W_{K}}\right\|_{F}=\mathcal{O}\left(\|W_{Q}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2}\right).$
For the attention with Pre-Norm and QK-Norm, we have
$\left\|\frac{\partial S}{\partial W_{O}}\right\|_{F}$ $=\mathcal{O}\left(\|W_{V}\|_{2}\right),\left\|\frac{\partial S}{\partial W_{V}}\right\|_{F}=\mathcal{O}\left(\|W_{O}\|_{F}\right),\left\|\frac{\partial S}{\partial W_{Q}}\right\|_{F}=\left\|\frac{\partial S}{\partial W_{K}}\right\|_{F}=\mathcal{O}\left(\|W_{V}\|_{2}\|W_{O}\|_{2}\right).$
For the attention with QKV-Norm, we have
$\left\|\frac{\partial S}{\partial W_{O}}\right\|_{F}$ $=\mathcal{O}\left(1\right),\left\|\frac{\partial S}{\partial W_{V}}\right\|_{F}=\mathcal{O}\left(\|W_{O}\|_{2}\right),\left\|\frac{\partial S}{\partial W_{Q}}\right\|_{F}=\left\|\frac{\partial S}{\partial W_{K}}\right\|_{F}=\mathcal{O}\left(\|W_{O}\|_{2}\right).$
The above theorem is an informal version of Theorem 2, with a more precise statement and proof provided in Appendix 8. In attention with Pre-Norm, gradients of weights exhibit strong dependencies on other weights; for instance, $W_{Q}$ and $W_{K}$ are influenced by all three other weights but not by themselves. In contrast, with QKV-Norm, the gradient of each weight depends at most on itself and $W_{O}$. This indicates that the gradient in Pre-Norm is more tightly coupled with other weights compared to QKV-Norm, while Pre-Norm with QK-Norm lies between the two. Consequently, during training, if the norm of a certain weight becomes excessively large, it is harder to control in Pre-Norm, leading to increased gradient magnitude. This creates a vicious cycle that may cause model collapse. In contrast, QKV-Norm alleviates this issue and significantly improves training stability. In summary, the degree of gradient coupling follows: Pre-Norm $>$ Pre-Norm with QK-Norm $>$ QKV-Norm; whereas training stability follows the reverse: Pre-Norm $<$ Pre-Norm with QK-Norm $<$ QKV-Norm.
5 Experiments
5.1 Experiment Settings
Baseline
We evaluate HybridNorm across two series of models: dense models and Mixture of Experts (MoE) models. The dense models include two scales: 550M and 1B, with the latter containing approximately 1.27 billion parameters and utilizing an architecture similar to LLaMA 3.2 [11]. All analytical experiments are conducted on the 550M dense models. For the MoE model, we use the OLMoE framework [24], which activates 1.3B parameters out of a total of 6.9B parameters (MoE-1B-7B). Both models are trained from scratch on the OLMoE Mix dataset [24].
Model Configuration
The 550M dense model has a model dimension of 1536, an FFN dimension of 4096, and utilizes 16 attention heads with 4 key/value heads per attention head. The 1.2B model features a larger model dimension of 2048 and an FFN dimension of 9192, with 32 attention heads and 8 key/value heads per attention head. The MoE-1B-7B model employs 16 attention heads, a model dimension of 2048, and an FFN dimension of 1024. Notably, it features 8 experts out of 64, providing a more fine-grained distribution of computational resources. All models consist of 16 layers and are trained with a consistent context length of 4096. More details can be found in Appendix 9.
Hyperparameters
Model weights are initialized using Megatron initialization [33] (See Section 5.3 for more details). For the optimization, we apply the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. All models are trained on sequences of 4096 tokens. For the dense model, we set the learning rate to 3e-4, decaying to 3e-5 using a cosine scheduler. The MoE model starts with a learning rate of 4e-4, decaying with a cosine schedule. We summarize the hyperparameters in Table 7.
Evaluation Metrics
To evaluate the performance of LLMs with HybridNorm, we employ a diverse set of open benchmarks, including ARC-Easy (ARC-E) [5], ARC-Challenge (ARC-C) [5], HellaSwag [46], PIQA [2], SciQ [41], CoQA [29], Winogrande [31], MMLU [15], BoolQ [4], COPA [13], CSQA [36], OBQA [23], and SocialIQA [32]. We leverage the LM Eval Harness [12] for standardized performance evaluation.
Figure 3: Training dynamics for 1.2B dense models with Pre-Norm, HybridNorm and HybridNorm under 1T training tokens. We present the training loss, validation loss, and downstream performance on HellaSwag and ARC-Easy, demonstrating that HybridNorm achieves superior performance.
Table 1: Downstream evaluation results of 1.2B dense models with Pre-Norm, HybridNorm, and HybridNorm under 1T training tokens. OQA refers to OpenbookQA.
Methods BasicArithmetic HellaSwag SciQ ARC-C ARC-E PIQA OQA COPA Avg.$\uparrow$
Pre-Norm 44.10 63.41 91.80 39.20 69.82 75.19 38.40 82.00 62.99
HybridNorm 44.12 64.22 91.88 39.13 71.05 74.72 38.88 82.00 63.25
HybridNorm 47.21 65.12 91.38 37.06 71.79 75.72 39.16 85.78 64.15
5.2 Main Results
Dense Models
We evaluate the performance of HybridNorm and HybridNorm on 1.2B dense transformer models. Figure 3 compares the training dynamics of dense models with different normalization methods. As shown in the figure, models with HybridNorm and HybridNorm exhibit consistently lower training loss and validation perplexity throughout training compared to Pre-Norm, highlighting their effectiveness in enhancing training stability and convergence. Table 1 presents the downstream evaluation results. HybridNorm consistently outperforms Pre-Norm in most tasks, achieving the highest average score. Notably, it demonstrates substantial improvements in tasks such as BasicArithmetic (+3.11), HellaSwag (+1.71), and COPA (+3.78), indicating enhanced generalization and robustness. These results underscore the scalability of HybridNorm in larger transformer models, further validating its effectiveness in improving both training stability and downstream performance. More results can be found in Figure 9. In Appendix 10.1, we present additional comparisons with other approaches, such as Post-Norm and Mix-LN. We further provide analyses of signal propagation and entropy dynamics across different methods, as detailed in Appendix 10.2 & 10.3.
MoE Models
For MoE models, we conduct experiments on MoE-1B-7B with 8 experts selected from a pool of 64. Figure 4 presents the training dynamics of MoE models under different normalization strategies. Throughout the training, HybridNorm consistently achieves lower training loss and validation perplexity compared to Pre-Norm. These findings indicate that HybridNorm effectively alleviates optimization difficulties in large-scale MoE models, resulting in more stable training and enhanced downstream performance. Further, as shown in Table 2, HybridNorm consistently outperforms Pre-Norm across various downstream tasks, achieving the highest average score. Notably, it demonstrates significant improvements in ARC-C (+2.35), ARC-E (+2.40), and OpenbookQA (+0.81), highlighting its ability to enhance generalization across diverse benchmarks.
Figure 4: Training dynamics for MoE-1B-7B models with Pre-Norm and HybridNorm under 500B training tokens. We present the training loss, validation loss, and downstream performance on HellaSwag and MMLU Var, demonstrating that HybridNorm achieves superior performance.
Table 2: Downstream evaluation results of MoE-1B-7B with Pre-Norm and HybridNorm under 500B training tokens. OQA refers to OpenbookQA.
Methods HellaSwag ARC-C ARC-E PIQA WinoGrande OQA BoolQ COPA Avg.$\uparrow$
Pre-Norm 69.94 39.92 73.37 77.82 63.34 42.37 67.47 85.40 64.95
HybridNorm 70.71 42.27 75.77 78.06 64.58 43.18 68.41 86.00 66.12
5.3 Ablation Studies
Table 3: Training loss and validation perplexity of 550M dense models with Pre-Norm and HybridNorm under various initialization methods and 400B training tokens.
[…]
Pre-Norm Normal 2.75 20.29 HybridNorm Normal 2.76 20.44
Depth-Scaled 2.76 20.49 Depth-Scaled 2.76 20.40
Megatron 2.76 20.44 Megatron 2.74 20.00
Initialization
To evaluate the sensitivity of Pre-Norm and HybridNorm to initialization schemes, we conduct ablation studies comparing three widely used initialization strategies: Normal initialization [25], Depth-Scaled initialization [48, 14], and Megatron initialization [33]. Normal initialization initializes all weights of linear layers using a truncated normal distribution with mean zero and standard deviation $1/\sqrt{2.5d}$, where $d$ is the hidden dimension. Depth-Scaled initialization and Megatron initialization introduce scaling factors to stabilize training in deep architectures. Specifically, Depth-Scaled initialization scales down the output projections of the attention and FFN by a factor of $\sqrt{2l}$, where $l$ is the layer index. In contrast, Megatron initialization scales down these projections by $\sqrt{2L}$, where $L$ is the total number of layers, mitigating gradient variance accumulation in very deep transformers. As shown in Table 3, Pre-Norm and HybridNorm exhibit sensitivity across different initialization methods, achieving the lowest training loss and perplexity under Normal initialization and Megatron initialization, respectively. Therefore, we set the default initialization method for Pre-Norm to Normal initialization and for HybridNorm to Megatron initialization in all experiments, respectively, which ensures that even under settings that may be more favorable to baseline models, the superiority of our approach is effectively demonstrated.
Table 4: Abalation study of the position of normalization layers on 550M dense models with 400B tokens. We report the training loss, the perplexity on C4 and Pile, and the accuracy on HS (HellaSwag).
[…]
QKVC-Post 2.74 20.05 10.34 52.68
QKC-Post 2.73 20.00 10.31 52.26
QK-Post - diverge
KV-Post 2.74 20.11 10.38 52.10
KC-Post 2.75 20.34 10.47 51.15
Pre-QKV-Post 2.74 20.13 10.37 52.65
Pre-QKV-Pre 2.74 19.97 10.33 53.05
Pre-QK-Pre 2.75 20.22 10.43 52.29
QKV-Pre 2.74 19.96 10.33 52.57
[…]
Post-Norm 2.76 20.43 10.57 51.20
Pre-Norm 2.75 20.30 10.48 51.97
QK-norm 2.75 20.22 10.43 52.29
Mix-LN 2.76 20.43 10.56 51.29
Post-Pre 2.75 20.26 10.46 51.19
Pre-Post 2.74 20.15 10.40 52.42
HybridNorm 2.74 20.00 10.29 53.35
HybridNorm 2.73 19.85 10.25 53.36
Normalization Position
We investigate the impact of the position of normalization layers within the transformer block. First, we examine the effect of varying the placement of QKV normalization (e.g., normalization setting in attention). We extend the normalization setting by considering not only the Query (Q), Key (K), and Value (V) components but also the Context (C), which refers to the output of the attention mechanism. For instance, QKVC-Norm applies normalization to all four components: Query, Key, Value, and Context, while KV-Norm and KC-Norm focus on the normalization of the Key-Value and Key-Context pairs, respectively. QKVC-Post refers to transformer blocks that employ QKVC-Norm in the MHA while using Post-Norm in the FFN. Second, we explore the effect of integrating QKV-Norm into different transformer architectures. For instance, Pre-QKV-Post refers to a configuration where QKV-Norm is applied with Pre-Norm in the MHA layer, while the FFN layer utilizes Post-Norm. Other configurations follow similar definitions. Finally, we compare various hybrid combinations of Pre-Norm and Post-Norm. Pre-Post refers to transformer blocks that apply Pre-Norm in the MHA and Post-Norm in the FFN, whereas Post-Pre adopts the opposite configuration. Mathematical formulas for methods mentioned above can be found in Appendix 12.
As shown in Table 4, HybridNorm (a.k.a. QKV-Post) and its variant HybridNorm consistently surpass other methods. Notably, HybridNorm achieves the lowest training loss and perplexity while attaining the highest accuracy on HellaSwag. Specifically, by comparing HybridNorm with the left first block in Table 4, we find that QKV-Norm is the most effective normalization. Similarly, comparing HybridNorm with the left second block, we observe that combining QKV-Norm with Post-Norm in the FFN yields superior performance. From the right table, one can see that the Pre-Post configuration indeed leads to improved performance, while replacing Pre-Norm in the MHA with QKV-Norm to form HybridNorm further enhances performance, achieving the best results.
Figure 5: Training loss and accuracy on HellaSwag of 550M dense models with different normalization methods for the first block.
Figure 6: Scaling law curves of Pre-Norm and HybridNorm.
Special Treatment of First Block
For the special treatment of the first block, we test different architectures, such as adding a normalization layer after embedding (call EmbedNorm) and armed the first block with QKV-norm and Pre-Norm in FFN (call First-QKV-Pre), which formulations are:
$Y^{0}$ $=\mathrm{MHA}_{QKV}(X^{0})+X^{0},\quad X^{1}=\mathrm{FFN}(\mathrm{Norm}(Y^{0}))+Y^{0}.$ (10)
As shown in Figure 6, we can see that, except for EmbedNorm, the special treatment of the first block effectively reduces training loss and improves downstream performance.
Table 5: Performance of dense models with different depths under 400B training tokens. We report the training loss, the perplexity on C4 and Pile, and the accuracy on HellaSwag and PIQA.
Models 550M, 16 Layers 543M, 29 Layers
[…]
Post-Norm 2.76 20.43 10.57 51.20 71.80 diverge
Pre-Norm 2.75 20.30 10.48 51.97 71.14 2.73 19.88 10.31 53.86 71.63
HybridNorm 2.74 20.00 10.29 53.35 71.96 2.72 19.67 10.18 54.89 72.62
HybridNorm 2.73 19.85 10.25 53.36 71.15 2.71 19.52 10.10 54.54 72.01
5.4 Scaling Laws Experiments
We compare the loss scaling curves between Pre-Norm and HybridNorm across a range of dense model sizes, from 151M to 1.2B parameters. The model sizes used for the scaling law experiments are detailed in Table 6, and all models are trained using the same setting and hyperparameters for fair comparison, as specified in Table 7. Models with 151M, 285M, 550M, and 1.2B parameters are trained on 200B, 200B, 300B, and 1T tokens, respectively. As shown in Figure 6, HybridNorm exhibits superior scaling properties, demonstrating lower training loss as the model size increases. This highlights its capacity to maintain both training stability and performance, even for extremely large models, thereby making it highly suitable for scaling to billion-parameter regimes.
5.5 Deeper Models
To further evaluate the robustness of HybridNorm and HybridNorm in deeper architectures, we conduct experiments on transformers with depths ranging from 16 to 29 layers while maintaining a comparable parameter budget. This setup allows for a fair comparison of different normalization strategies in deep transformer architectures. As shown in Table 5, both HybridNorm and HybridNorm consistently outperform Pre-Norm and Post-Norm across various depths, demonstrating their effectiveness in stabilizing deep model training. A particularly striking observation is that Post-Norm fails to converge at 29 layers, reinforcing its well-documented instability in deeper architectures. In contrast, HybridNorm and HybridNorm not only ensure stable training across all depths but also achieve significantly lower training loss and perplexity on both the C4 and Pile datasets. These improvements indicate that HybridNorm-based normalization strategies mitigate optimization difficulties that commonly arise in deep transformers. Furthermore, HybridNorm achieves the highest accuracy on challenging downstream benchmarks such as HellaSwag and PIQA, suggesting that its benefits extend beyond mere training stability to enhanced generalization on real-world tasks. These results provide strong empirical evidence that HybridNorm-based normalization schemes enable deeper transformer training while preserving superior optimization efficiency and downstream task performance.
6 Conclusion
In this paper, we have presented HybridNorm, a novel hybrid normalization approach that effectively integrates the strengths of Pre-Norm and Post-Norm to address the longstanding trade-offs in transformer training. We have provided both theoretical and empirical insights into how HybridNorm stabilizes gradient flow while maintaining strong regularization effects, demonstrating its potential to improve both convergence and final model performance. Extensive experiments across diverse benchmarks validate the efficacy of our approach, with results consistently showing that HybridNorm outperforms traditional normalization strategies in terms of stability and accuracy. Our findings underscore the necessity of re-evaluating normalization position within transformer architectures, thereby laying the groundwork for further exploration of hybrid strategies. We posit that HybridNorm signifies a substantial advancement in the design of more robust and performant deep transformer models, offering practical advantages for the training of next-generation large-scale neural networks.